# Mixing Classifiers to Alleviate the Accuracy-Robustness Trade-Off

### v1

@@ Abstract
Deep learning classifiers have recently found tremendous success in data-driven control systems. However, standard learning models often suffer from an accuracy-robustness trade-off, which is a limitation that must be overcome in the control of safety-critical systems that require both high performance and rigorous robustness guarantees. In this work, we build upon the recent “locally biased smoothing” method to develop classifiers that simultaneously inherit high accuracy from standard models and high robustness from robust models. Specifically, we extend locally biased smoothing to the multi-class setting, and then overcome its performance bottleneck by generalizing the formulation to “mix” the outputs of a standard neural network and a robust neural network. Both models are pre-trained, and thus our method does not require additional training. We prove that when the robustness of the robust base model is certifiable, no alteration or attack within a closed-form $\ell_{p}$ radius on an input can result in misclassification of the mixed classifier; the proposed model inherits the certified robustness. Moreover, we use numerical experiments on the CIFAR-10 benchmark dataset to verify that the mixed model noticeably improves the accuracy-robustness trade-off.
I Introduction
@@
In recent years, high-performing machine learning models have been successfully employed in a range of control settings, including reinforcement learning for dynamical systems with uncertainty [1, 2] and self-driving cars [3, 4]. However, models such as neural networks have been shown to be vulnerable to adversarial attacks, which are imperceptibly small input data alterations maliciously designed to cause failure [5, 6, 7]. For example, both digital and physical attacks on traffic signs have successfully fooled state-of-the-art image classifiers [8, 9]. This vulnerability makes such models unreliable for safety-critical control where guaranteeing robustness is necessary. In response, “adversarial training (AT)” [10, 11, 12, 13, 14] has been studied to alleviate the susceptibility. AT builds robust neural networks by training on adversarially attacked data.
@@
A parallel line of work focuses on certified (that is, mathematical proof of) robustness [15, 16, 17]. Among the most popular of these methods, “randomized smoothing (RS)” seeks to achieve certified robustness by processing intentionally corrupted data at test time [18, 19, 20]. RS has recently been applied to robustify reinforcement learning-based control strategies [21, 22]. The recent work [23] has shown that a “locally biased smoothing” method, which robustifies the model locally based on the particular test datum, outperforms the traditional data-agnostic RS that uses globally fixed smoothing noise. However, [23] only focuses on binary classification problems, significantly limiting the applications. Moreover, [23] relies on the robustness of a $K$-nearest-neighbor ($K$-NN) classifier, which suffers from the lack of representation power when applied to harder problems and becomes a performance bottleneck.
@@
While some works have shown that there exists a fundamental trade-off between accuracy and robustness [24, 25], recent research has argued that it should be possible to simultaneously achieve robustness and accuracy on benchmark datasets [26]. To this end, variants of AT that improve the accuracy-robustness trade-off have been proposed, including TRADES [25], Interpolated Adversarial Training [27], and many others [28, 29, 30, 31]. However, even with these improvements, degraded clean accuracy is often an inevitable price of achieving robustness. Moreover, standard non-robust models often achieve enormous performance gains by pre-training on larger datasets, whereas the effect of pre-training on robust classifiers is less understood and may be less prominent [32, 33].
@@
This work makes a theoretically disciplined step towards robustifying models without sacrificing clean accuracy. Specifically, we build upon locally biased smoothing and replace its underlying $K$-NN classifier with a robust neural network that can be obtained via various existing methods. We also modify how the standard and robust models are “mixed” accordingly. The resulting formulation, to be introduced formally in Section III, is a convex combination of the output of a standard neural network and the output of a robust neural network. We prove that, when the robust neural network has a bounded Lipschitz constant or is built via RS, the mixed classifier also has a closed-form certified robust radius. More importantly, the proposed method achieves an empirical robustness level close to that of the robust base model while approaching the clean accuracy of the standard base classifier. This desirable behavior significantly improves the accuracy-robustness trade-off for tasks where standard models noticeably outperform robust models on clean data.
@@
Note that we do not make any assumptions about how the standard and robust base models are obtained (can be AT, RS, or others), nor do we make assumptions on the adversarial attack type and budget. Thus, our mixed classification scheme can take advantage of pre-training on large datasets via the standard base classifier and benefit from ever-improving robust training methods via the robust base classifier.
II Background and related works
II-A Notations
@@
II-B Related Adversarial Attacks and Defenses
The fast gradient sign method (FGSM) and projected gradient descent (PGD) attacks based on differentiating the cross-entropy loss have been considered the most classic and straightforward attacks [34, 11]. However, they have been shown to be too weak, as defenses that are only designed against these attacks can be easily circumvented [35, 36, 37]. To this end, various attack methods based on alternative loss functions, Expectation Over Transformation, and black-box perturbations have been proposed. Such efforts include AutoAttack [38], adaptive attack [39], and many others.
@@ II-B Related Adversarial Attacks and Defenses
On the defense side, while AT [34] and TRADES [25] have seen enormous success, such methods are often limited by a significantly larger amount of required training data [40] and a decrease in generalization capability. Initiatives that construct more effective training data via data augmentation [41, 42] and generative models [43] have successfully produced more robust models. Improved versions of AT [44, 45] have also been proposed.
@@
Previous initiatives that aim to enhance the accuracy-robustness trade-off include using alternative attacks during training [46], appending early-exit side branches to a single network [47], and applying AT for regularization [48]. Moreover, ensemble-based defenses, such as random ensemble [49] and diverse ensemble [50, 51], have been proposed. In comparison, this work considers two separate classifiers and uses their synergy to improve the accuracy-robustness trade-off, achieving higher performances.
II-C Locally Biased Smoothing
@@
Randomized smoothing, popularized by [18], achieves robustness at test time by replacing $f(x)=\operatorname*{arg\,max}_{i\in[c]}g_{i}(x)$ with a smoothed classifier $\widetilde{f}(x)=\operatorname*{arg\,max}_{i\in[c]}\mathbb{E}_{\xi\sim{\mathcal{S}}}\left[g(x+\xi)\right]$, where ${\mathcal{S}}$ is a smoothing distribution. A common choice for ${\mathcal{S}}$ is a Gaussian distribution.
@@
The authors of [23] have recently argued that data-invariant RS does not always achieve robustness. They have shown that in the binary classification setting, RS with an unbiased distribution is suboptimal, and an optimal smoothing procedure shifts the input point in the direction of its true class. Since the true class is generally unavailable, a “direction oracle” is used as a surrogate. This “locally biased smoothing” method is no longer randomized and outperforms traditional data-blind RS. The locally biased smoothed classifier, denoted $h^{\gamma}\colon{\mathbb{R}}^{d}\to{\mathbb{R}}$, is obtained via the deterministic calculation
$h^{\gamma}(x)=g(x)+\gamma h(x)\lVert\nabla g(x)\rVert_{p*},$
where $h(x)\in\{-1,1\}$ is the direction oracle and $\gamma\geq 0$ is a trade-off parameter. The direction oracle should come from an inherently robust classifier (which is often less accurate). In [23], this direction oracle is chosen to be a one-nearest-neighbor classifier.
III Using a Robust Neural Network as the Smoothing Oracle
@@
Since locally biased smoothing was designed for binary classification problems, we first extend it to the multi-class setting. To achieve this, we treat the output $h^{\gamma}_{i}(x)$ of each class independently, giving rise to:
@@
Note that if $\lVert\nabla g_{i}(x)\rVert_{p*}$ is large for some $i$, then $h_{\text{smo1},i}^{\gamma}(x)$ can be large even if both $g_{i}(x)$ and $h_{i}(x)$ are small, potentially leading to incorrect predictions. To remove the effect of the magnitude difference across the classes, we propose a normalized formulation as follows:
@@
The parameter $\gamma$ adjusts the trade-off between clean accuracy and robustness. When $\gamma=0$, it holds that $h_{\text{smo2},i}^{\gamma}(x)\equiv g_{i}(x)$ for all $i$. When $\gamma\to\infty$, it holds that $h_{\text{smo2},i}^{\gamma}(x)\to h_{i}(x)$ for all $x$ and all $i$.
With the mixing procedure generalized to the multi-class setting, we now discuss the choice of the smoothing oracle $h_{i}(\cdot)$. While $K$-NN classifiers are relatively robust and can be used as the oracle, their representation power is too weak. On the CIFAR-10 image classification task [52], $K$-NN only achieves around $35\%$ accuracy on clean test data. In contrast, an adversarially trained ResNet can reach $50\%$ accuracy on attacked test data [34]. This lackluster performance of $K$-NN becomes a significant bottleneck in the accuracy-robustness trade-off of the mixed classifier. To this end, we replace the $K$-NN model with a robust neural network. The robustness of this network can be achieved via various methods, including AT, TRADES, and RS.
Further scrutinizing Eq. 2 leads to the question of whether $\lVert\nabla g_{i}(x)\rVert_{p*}$ is the best choice for adjusting the mixture of $g(\cdot)$ and $h(\cdot)$. In fact, this gradient magnitude term is a result of the assumption that $h(x)\in\{-1,1\}$, which is the setting considered in [23]. Here, we no longer have this assumption. Instead, we assume both $g(\cdot)$ and $h(\cdot)$ to be differentiable. Thus, we further generalize the formulation to
@@
Another design question is whether $g(\cdot)$ and $h(\cdot)$ should be the pre-softmax logits or the post-softmax probabilities. Note that since most attack methods are designed based on the logits, incorporating the softmax function into the model may result in gradient masking, an undesirable phenomenon that makes it hard to properly evaluate the proposed method. Thus, we have the following two options that make the mixed model compatible with existing gradient-based attacks:
@@
    Use the logits for both $g(\cdot)$ and $h(\cdot)$.

### v2

@@ Abstract
Deep neural classifiers have recently found tremendous success in data-driven control systems. However, existing models suffer from a trade-off between accuracy and adversarial robustness. This limitation must be overcome in the control of safety-critical systems that require both high performance and rigorous robustness guarantees. In this work, we develop classifiers that simultaneously inherit high robustness from robust models and high accuracy from standard models. Specifically, we propose a theoretically motivated formulation that mixes the output probabilities of a standard neural network and a robust neural network. Both base classifiers are pre-trained, and thus our method does not require additional training. Our numerical experiments verify that the mixed classifier noticeably improves the accuracy-robustness trade-off and identify the confidence property of the robust base classifier as the key leverage of this more benign trade-off. Our theoretical results prove that under mild assumptions, when the robustness of the robust base model is certifiable, no alteration or attack within a closed-form $\ell_{p}$ radius on an input can result in misclassification of the mixed classifier.
keywords:
Adversarial Robustness, Image Classification, Computer Vision, Model Ensemble
1 Introduction
@@
In recent years, high-performance machine learning models have been employed in various control settings, including reinforcement learning for dynamic systems with uncertainty (Levine et al., 2016; Sutton and Barto, 2018) and autonomous driving (Bojarski et al., 2016; Wu et al., 2017). However, models such as neural networks have been shown to be vulnerable to adversarial attacks, which are imperceptibly small input data alterations maliciously designed to cause failure (Szegedy et al., 2014; Nguyen et al., 2015; Huang et al., 2017; Eykholt et al., 2018; Liu et al., 2019). This vulnerability makes such models unreliable for safety-critical control where guaranteeing robustness is necessary. In response, “adversarial training (AT)” (Kurakin et al., 2017; Goodfellow et al., 2015; Bai et al., 2022a, b; Zheng et al., 2020; Zhang et al., 2019) have been studied to alleviate the susceptibility. AT builds robust neural networks by training on adversarially attacked data.
@@
A parallel line of work focuses on mathematically certified robustness (Anderson et al., 2020; Ma and Sojoudi, 2021; Anderson and Sojoudi, 2022a). Among these methods, “randomized smoothing (RS)” is a particularly popular one that seeks to achieve certified robustness by processing intentionally corrupted data at inference time (Cohen et al., 2019; Li et al., 2019; Pfrommer et al., 2023), and has recently been applied to robustify reinforcement learning-based control strategies (Kumar et al., 2022; Wu et al., 2022). The recent work (Anderson and Sojoudi, 2022b) has shown that “locally biased smoothing,” which robustifies the model locally based on the input test datum, outperforms the traditional RS with fixed smoothing noise. However, Anderson and Sojoudi (2022b) only focus on binary classification problems, significantly limiting the applications. Moreover, Anderson and Sojoudi (2022b) rely on the robustness of a $K$-nearest-neighbor ($K$-NN) classifier, which suffers from a lack of representation power when applied to harder problems and becomes a bottleneck.
@@
While some works have shown that there exists a fundamental trade-off between accuracy and robustness (Tsipras et al., 2019; Zhang et al., 2019), recent research has argued that it should be possible to simultaneously achieve robustness and accuracy on benchmark datasets (Yang et al., 2020). To this end, variants of AT that improve the accuracy-robustness trade-off have been proposed, including TRADES (Zhang et al., 2019), Interpolated Adversarial Training (Lamb et al., 2019), and many others (Raghunathan et al., 2020; Zhang and Wang, 2019; Tramèr et al., 2018; Balaji et al., 2019). However, even with these improvements, degraded clean accuracy is often an inevitable price of achieving robustness. Moreover, standard non-robust models often achieve enormous performance gains by pre-training on larger datasets, whereas the effect of pre-training on robust classifiers is less understood and may be less prominent (Chen et al., 2020; Fan et al., 2021).
@@
This work makes a theoretically disciplined step towards robustifying models without sacrificing clean accuracy. Specifically, we build upon locally biased smoothing and replace its underlying $K$-NN classifier with a robust neural network that can be obtained via various existing methods. We then modify how the standard base model (a highly accurate but possibly non-robust neural network) and the robust base model are “mixed” accordingly. The resulting formulation, to be introduced in Section 3, is a convex combination of the output probabilities from the two base classifiers. We prove that, when the robust network has a bounded Lipschitz constant or is built via RS, the mixed classifier also has a closed-form certified robust radius. More importantly, our method achieves an empirical robustness level close to that of the robust base model while approaching the standard base model’s clean accuracy. This desirable behavior significantly improves the accuracy-robustness trade-off, especially for tasks where standard models noticeably outperform robust models on clean data.
@@
Note that we do not make any assumptions about how the standard and robust base models are obtained (can be AT, RS, or others), nor do we assume the adversarial attack type and budget. Thus, our mixed classification scheme can take advantage of pre-training on large datasets via the standard base classifier and benefit from ever-improving robust training methods via the robust base classifier.
2 Background and related works
2.1 Notations
@@
2.2 Related Adversarial Attacks and Defenses
The fast gradient sign method (FGSM) and projected gradient descent (PGD) attacks based on differentiating the cross-entropy loss are highly effective and have been considered the most standard attacks for evaluating robust models (Madry et al., 2018; Goodfellow et al., 2015). To exploit the structures of the defense methods, adaptive attacks have also been introduced (Tramèr et al., 2020).
@@ 2.2 Related Adversarial Attacks and Defenses
On the defense side, while AT (Madry et al., 2018) and TRADES (Zhang et al., 2019) have seen enormous success, such methods are often limited by a significantly larger amount of required training data (Schmidt et al., 2018) and a decrease in generalization capability. Initiatives that construct more effective training data via data augmentation (Rebuffi et al., 2021; Gowal et al., 2021) and generative models (Sehwag et al., 2022) have successfully produced more robust models. Improved versions of AT (Jia et al., 2022; Shafahi et al., 2019) have also been proposed.
@@
Previous initiatives that aim to enhance the accuracy-robustness trade-off include using alternative attacks during training (Pang et al., 2022), appending early-exit side branches to a single network (Hu et al., 2020), and applying AT for regularization (Zheng et al., 2021). Moreover, ensemble-based defenses, such as random ensemble (Liu et al., 2018) and diverse ensemble (Pang et al., 2019; Alam et al., 2022), have been proposed. In comparison, this work considers two separate classifiers and uses their synergy to improve the accuracy-robustness trade-off, achieving higher performances.
2.3 Locally Biased Smoothing
@@
Randomized smoothing, popularized by (Cohen et al., 2019), achieves robustness at inference time by replacing $f(x)=\operatorname*{arg\,max}_{i\in[c]}g_{i}(x)$ with a smoothed classifier $\widetilde{f}(x)=\operatorname*{arg\,max}_{i\in[c]}\mathbb{E}_{\xi\sim{\mathcal{S}}}\left[g_{i}(x+\xi)\right]$ , where ${\mathcal{S}}$ is a smoothing distribution. A common choice for ${\mathcal{S}}$ is a Gaussian distribution.
@@
Anderson and Sojoudi (2022b) have recently argued that data-invariant RS does not always achieve robustness. They have shown that in the binary classification setting, RS with an unbiased distribution is suboptimal, and an optimal smoothing procedure shifts the input point in the direction of its true class. Since the true class is generally unavailable, a “direction oracle” is used as a surrogate. This “locally biased smoothing” method is no longer randomized and outperforms traditional data-blind RS. The locally biased smoothed classifier, denoted $h^{\gamma}\colon{\mathbb{R}}^{d}\to{\mathbb{R}}$, is obtained via the deterministic calculation $h^{\gamma}(x)=g(x)+\gamma h(x)\lVert\nabla g(x)\rVert_{p*}$, where $h(x)\in\{-1,1\}$ is the direction oracle and $\gamma\geq 0$ is a trade-off parameter. The direction oracle should come from an inherently robust classifier (which is often less accurate). In (Anderson and Sojoudi, 2022b), this direction oracle is chosen to be a one-nearest-neighbor classifier.
3 Using a Robust Neural Network as the Smoothing Oracle
@@
Locally biased smoothing was designed for binary classification, restricting its practicality. Here, we first extend it to the multi-class setting by treating the output of each class, denoted as $h^{\gamma}_{i}(x)$, independently, giving rise to:
@@
Note that if $\lVert\nabla g_{i}(x)\rVert_{p*}$ is large for some class $i$, then $h_{\text{smo1},i}^{\gamma}(x)$ can be large for class $i$ even if both $g_{i}(x)$ and $h_{i}(x)$ are small, leading to incorrect predictions. To remove the effect of the gradient magnitude difference across the classes, we propose a normalized formulation as follows:
@@
The parameter $\gamma$ adjusts between clean accuracy and robustness. It holds that $h_{\text{smo2},i}^{\gamma}(x)\equiv g_{i}(x)$ when $\gamma=0$, and $h_{\text{smo2},i}^{\gamma}(x)\to h_{i}(x)$ when $\gamma\to\infty$ for all $x$ and all $i$.
With the mixing procedure generalized to the multi-class setting, we now discuss the choice of the smoothing oracle $h_{i}(\cdot)$. While $K$-NN classifiers are relatively robust and can be used as the oracle, their representation power is too weak. On the CIFAR-10 image classification task (Krizhevsky, 2012), $K$-NN only achieves around $35\%$ accuracy on clean test data. In contrast, an adversarially trained ResNet can reach $50\%$ accuracy on attacked test data (Madry et al., 2018). This lackluster performance of $K$-NN becomes a significant bottleneck in the accuracy-robustness trade-off of the mixed classifier. To this end, we replace the $K$-NN model with a robust neural network. The robustness of this network can be achieved via various methods, including AT, TRADES, and RS.
Further scrutinizing Eq. 2 leads to the question of whether $\lVert\nabla g_{i}(x)\rVert_{p*}$ is the best choice for adjusting the mixture of $g(\cdot)$ and $h(\cdot)$. This gradient magnitude term is a result of Anderson and Sojoudi (2022b)’s assumption that $h(x)\in\{-1,1\}$. Here, we no longer have this assumption. Instead, we assume both $g(\cdot)$ and $h(\cdot)$ to be differentiable. Thus, we generalize the formulation to
@@
Another design question is whether $g(\cdot)$ and $h(\cdot)$ should be the pre-softmax logits or the post-softmax probabilities. Note that since most attack methods are designed based on logits, the output of the mixed classifier should be logits rather than probabilities to avoid gradient masking, an undesirable phenomenon that makes it hard to evaluate the robustness properly. Thus, we have the following two options that make the mixed model compatible with existing gradient-based attacks:
@@
    Use the logits for both base classifiers, $g(\cdot)$ and $h(\cdot)$.